A car is traveling due north at 22.2 m/s.

Part A

Find the velocity of the car after 5.07 s if its acceleration is 1.18 m/s2 due north

Find the velocity of the car after 5.07 s if its acceleration is 1.09 m/s2 due south
Answer #1

Solution) u = 22.2 m/s ( north )

North is upward represented by (+j)

South is downward represented by (-j)

u = 22.2(j)

(A) v = ?

t = 5.07 s

a = 1.18 m/s^2 (north)

a = 1.18(j)

v = u + at

v = 22.2(j) + 1.18(j)(5.07)

v = 28.18(j) m/s

* a = 1.09(- j) m/s^2

v = u + at

v = 22.2(j) + 1.09(-j)(5.07)

v = 16.67(j) m/s